What is the fourth term of the sequence whose first term is a1=2 and whose nth term is an=4a(n-1)-11
My problem here is that i can't figure out how to use the nth term, what do you do with the info given?

Answer by jim_thompson5910(35256)

This is a recursive sequence. This means that you need to find each term before the fourth term a4 to find a4.


First term: a1 = 2

Second term: a2 = 4*a1-11 = 4*2-11 = 8-11 = -3

Third term: a3 = 4*a2-11 = 4*(-3)-11 = -12-11 = -23

Fourth term: a4 = 4*a3-11 = 4*(-23)-11 = -92-11 = -103

So the fourth term is a4 = -103